OpenStudy (anonymous):

HELP ME PLEASE: A car is traveling on a highway. The distance (in miles) from its destination and the time (in hours) is given by the equation d=420-65t What is the practical meaning of the d-intercept? a. Initially the car is 355 miles from its destination. b. Initially the car is 420 miles from its destination. c. Initially the car is 65 miles from its destination. d. Initially the car is 485 miles from its destination.

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

PLEASE HELP ME !

OpenStudy (anonymous):

@dumbcow

OpenStudy (debbieg):

The d-intercept (which is really just what you are used to calling the y-intercept, here we are just using a different variable) occurs where t=0 (what you are used to seeing as x).

OpenStudy (debbieg):

So if t=0, what does that mean IN THE CONTEXT of this problem? What is that, in terms of the model?

OpenStudy (anonymous):

that d= 0 ?

OpenStudy (debbieg):

No, if t=0, what does d=? look at your equation: d=420-65t

OpenStudy (debbieg):

Remember, the d-intercept is the value of d, when t=0. What is that value, for THIS equation? And what time is it when t=0?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

because anything multiplied by zero = 0

OpenStudy (debbieg):

d=420-65t Put in t=0: d=420-65(0) Yes, anything multiplied by 0, =0. So when t=0, WHAT IS d?

OpenStudy (anonymous):

d=0

OpenStudy (anonymous):

0,420 ?

OpenStudy (debbieg):

When t=0 what does d=? I can't tell what you mean by 0,420.

OpenStudy (debbieg):

Just tell me: when t=0, what does d=?

OpenStudy (anonymous):

420 ?

OpenStudy (anonymous):

wait d = 65

OpenStudy (anonymous):

Im sorry im confused

OpenStudy (debbieg):

Yes!! OK... now we are getting somewhere. :) so when t=0, d=420. Now remember, that's the d-intercept, right?

OpenStudy (debbieg):

Oh crap, I thought you had it. lol. let's back up.

OpenStudy (anonymous):

Seee I did have it, and u got me confused lol :)

OpenStudy (debbieg):

d=420-65t Put in t=0: d=420-65(0) As you said, anything x 0 = 0, so you have: d=420-0=420 Got that?

OpenStudy (anonymous):

Yes I got tht..

OpenStudy (debbieg):

Nothing *I* said could have given you the idea that d=65 when t=0. ;) OK, so the question is, "What is the practical meaning of the d-intercept?" You now know that the d-intercept=420, because that's what you get for d (distance from destination) when t=0. Now remembering that t represents time, what does IT MEAN for t to be =0?

OpenStudy (anonymous):

I don't get what you're asking?

OpenStudy (debbieg):

t represents time. So when t=0, what time are you at "in the process" being modeled? When t=1, you are 1 hour in (1 hour has passed). When t=5, you are five hours in. where are you when t=0? How much time has passed?